# A Multimodal Deviation Perceiving Framework for Weakly-Supervised Temporal Forgery Localization

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33th ACM International Conference on Multimedia; October 27–31, 2025; Ireland, Dublinbooktitle: Proceedings of the 33th ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Ireland, Dublinccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer visionccs: Computing methodologies Computer vision problems
@@ 1. Introduction
Current research in Deepfake forensics primarily tackles the issue through classification tasks, particularly binary classification for videos or images (Yin et al., 2023; Gu et al., 2022; Dong et al., 2023). Nevertheless, this methodology exhibits limitations when addressing more challenging deepfake scenarios, particularly in the context of temporal partial forgery localization. Considering the specificity and potential pernicious effects of temporal partial forgery, Chugh  (2020) proposed the temporal forgery localization task (TFL) to localize the start and end timestamps of forged segments. Several researches (Cai et al., 2023; Liu et al., 2023) have explored the TFL task with a fully-supervised methodology. Both BA-TFD+ (Cai et al., 2023) and AVTFD (Liu et al., 2023) attempted to combine the TFL with frame-level Deepfake detection methods. UMMAFormer (Zhang et al., 2023) aimed to mine forgery traces through feature reconstruction. The aforementioned fully-supervised temporal forgery localization (FS-TFL) methods have achieved some degree of localization performance. However, they require elaborate frame-level or timestamp annotations for fully-supervised learning, which is usually costly and time-consuming.
@@ 3.4. Deviation Perceiving loss
where $f\left(x_{t},x_{t+1}\right)$ indicates the deviation between the $t$-th segment and the $(t+1)$-th segment, and $f(\cdot)$ is a deviation measure function that measures the degree of deviation (e.g., mean square error (MSE) as shown in Eq. (10)). We assessed the impact of different $f(\cdot)$ on the performance of temporal forgery localization in Section 4.5.
@@ 4.1. Experimental Setup
Datasets: We conduct experiments on two challenging Deepfake temporal partial forgery datasets LAV-DF (Cai et al., 2023) and AV-Deepfake1M (Cai et al., 2024). LAV-DF is a strategic content-driven multimodal forgery dataset, which contains $36,431$ genuine videos and $99,873$ forged videos. The duration of forged segments is in the range of $[0-1.6s]$. AV-Deepfake1M is a large-scale multimodal forgery dataset which contains $2,068$ subjects resulting in $286,721$ genuine videos and $860,039$ forged videos. There are four types of samples (real, visual-only forgery, audio-only forgery and audio-visual forgery) in both LAV-DF and AV-Deepfake1M.
@@
Evaluation metrics: The average precision (AP) and average recall (AR) are utilized as the evaluation metrics following (Cai et al., 2023) and (Zhang et al., 2023). For LAV-DF, the IoU thresholds of AP are set as $0.5$, $0.75$ and $0.95$. As AV-Deepfake1M is a more challenging dataset compared to LAV-DF, the IoU thresholds of AP are set as $[0.1:0.1:0.7]$. AR is calculated using $20$, $10$, $5$, and $2$ proposals with IoU thresholds $[0.5:0.05:0.95]$, respectively. Implementation details: The MDP is trained by Adam optimizer with a learning rate of $51e-$ 5, a batch size of 32. The hyperparameter $\phi$ is set as $0.5$.
@@ 4.2. Performance Comparisons
LAV-DF Dataset: As shown in Table 1, the results show that MDP, which is a weakly-supervised temporal forgery localization approach, achieves relatively good performance on both AP and AR. Compared to the weakly-supervised temporal action localization approaches, the MDP is significantly improved in both AP and AR. It could be found that the AR@$2$ to AR@$20$ remain consistent, which indicates the proposed framework could predict the forged segments with less number of candidate proposals. As for the AP evaluation metric, MDP achieves significantly superior performance compared to the comparison weakly-supervised approaches on both AP@$0.5$ and AP@$0.75$, which indicates that MDP exhibits higher accuracy in predicting forged segments. It should be noted that despite MDP showing satisfactory performance on most evaluation metrics, it underperforms on the AP@$0.95$, implying that it is inadequate in localizing the precise timestamps of the forged segments.
@@
Obviously, compared to the weakly-supervised approaches, the fully-supervised approaches achieve superior performance on both AP and AR. Such results are also reasonable, as fully-supervised approaches are more adept at learning the relationship between the Deepfake video features and corresponding timestamps of forged segments with provided frame-level annotations. Nevertheless, it could be observed that the MDP also achieves relatively good performance on AP@$0.5$ and AP@$0.75$ with a relatively small gap with fully-supervised approaches. The experimental results indicate that, despite utilizing only video-level annotations for weakly-supervised learning, the MDP could still effectively identify temporal forgery traces present within the multimodal features, enabling relatively precise localization of timestamps for forged segments. This is achieved by analyzing the temporal deviations in the multimodal features, demonstrating the MDP’s performance in addressing temporal forgery localization.
AV-Deepfake1M Dataset: Compared to LAV-DF, AV-Deepfake1M contains more Deepfake videos with long duration. The longest duration of Deepfake video in LAV-DF is $19.97$s, while the longest video in AV-Deepfake1M is $32.51$s, and $4\%$ samples of the AV-Deepfake1M have a duration of more than $20$s. Long duration videos are more challenging for weakly-supervised temporal forgery localization task that depend merely on video-level annotations. Since the comparison weakly-supervised approaches could not localize the timestamps of forged segments in AV-Deepfake1M effectively, we have chosen the fully-supervised approaches (ActionFormer, TriDet, UMMAFormer and MFMS) to compare with the MDP on this dataset.
@@
The experimental results are shown in Table 2. It could be observed that both fully-supervised approaches and weakly-supervised approach MDP have decreased in AR evaluation metric compared to the experimental results of LAV-DF. As for AP, the proposed MDP underperforms on AP evaluation metrics with large IoU thresholds, while it achieves localization results close to the fully-supervised UMMAFormer on AP@$0.1$ and AP@$0.2$. The experimental results show that MDP could mine temporal forgery traces and localize the timestamps of forged segments even in challenging AV-Deepfake1M dataset by relying merely on the video-level annotations. We also conducted experiments with WS-TAL approaches CoLA and FuSTAL on the AV-Deepfake1M dataset. The corresponding experimental results were not displayed in Table  2 since the obtained localization results are not effective. The comparison results on this dataset also illustrate that the MDP could exploit the subtle forgery traces within the temporal features, thereby addressing the task of timestamps localization of the forged segments relying on the video-level annotations.
@@
4.3. Ablation Study
It could be observed that the localization performance is significantly improved on both average AP ($+14.39\%$) and average AR ($+9.8\%$) with the enhancement of the cross-modal attention mechanism. In multimodal Deepfake scenarios, the visual and audio modalities are typically embedded with extensive features, which are critical for mining forgery traces. In the spatial domain, the visual modality contains richer information compared to the audio modality. Conversely, In the temporal domain, the audio modality possesses a higher frequency of sampling points per unit of time than the visual modality. Therefore, the proposed MDP aligns the visual and audio features in spatial and temporal domains. Both visual and audio features are transformed into 1-D feature vectors on the spatial domain. In the temporal domain, they are scaled to a uniform time dimension through a pooling operation. This ensures the effective integration of multimodal data for better analysis. For WS-TFL, the temporal information is apparently essential for the temporal forgery localization. Subsequently, the cross-modal attention is utilized to enhance the audio features and visual features in a temporal property preservation manner, respectively. The experimental results of the ablation study further validate the effectiveness of the proposed cross-modal attention component.
In addition, compared to the baseline, the introduction of deviation perceiving loss $\mathcal{L}_{dp}$ also improves the performance of temporal forgery localization on both average AP ($+11.71\%$) and average AR ($+1.89\%$). Because WS-TFL could merely utilize the video-level annotations for model parameter learning, there is no temporal information to guide the model training, and locating the timestamps of forged segments is challenging in this background. The previous Deepfake detection approaches have demonstrated that forged samples tend to have a larger deviation between adjacent segments compared to genuine samples. The $\mathcal{L}_{dp}$ is based on measuring the deviation between adjacent segments, which consequently constrains the MDP to enlarge the temporal deviation of forgery samples while reducing that of genuine samples. The results of the ablation study indicate that $\mathcal{L}_{dp}$ could guide the MDP in perceiving the deviation between adjacent segments, and thus mine more temporal information for temporal forgery localization. According to Table 3, the best average AP and average AR are achieved by introducing both CMA and $\mathcal{L}_{dp}$. It validates the effectiveness of the key components in MDP.
4.4. Visualization Analysis
From the visualization results, it could be observed that MDP could localize the timestamps of all the forged segments more precisely compared to the CoLA and FuSTAL on LAV-DF. Moreover, the localization effectiveness of MDP is comparable to that observed in fully-supervised approaches within the presented sample. In particular, the AV-Deepfake1M sample presents a notable challenge, as the duration of all three forged segments is below $0.3$s, while the overall duration of the Deepfake video is $23.36$s. The localization of a small proportion of temporal forgeries is a significant challenge for WS-TFL. Despite this, the MDP effectively utilizes weakly-supervised learning based solely on video-level annotations. It accurately predicts the timestamps of the forged segments, achieving comparable localization results to that obtained through fully-supervised approaches training at the frame-level annotations. This represents that the proposed MDP effectively leverages the temporal forgery traces following the interaction of multimodal features, thereby enabling the precise identification of both the start and end timestamps of the forged segments.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755534isbn: 979-8-4007-2035-2/2025/10submissionid: 4314ccs: Computing methodologiesccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer visionccs: Computing methodologies Computer vision problems
@@ 1. Introduction
Current research in Deepfake forensics primarily tackles the issue through classification tasks, particularly binary classification for videos or images (Yin et al., 2023; Gu et al., 2022; Dong et al., 2023). Nevertheless, this methodology exhibits limitations when addressing more challenging deepfake scenarios, particularly in the context of temporal partial forgery localization. Considering the specificity and potential pernicious effects of temporal partial forgery, Chugh  (2020) proposed the temporal forgery localization task (TFL) to localize the start and end timestamps of forged segments. Several researches (Cai et al., 2023; Liu et al., 2024) have explored the TFL task with a fully-supervised methodology. Both BA-TFD+ (Cai et al., 2023) and AVTFD (Liu et al., 2024) attempted to combine the TFL with frame-level Deepfake detection methods. UMMAFormer (Zhang et al., 2023) aimed to mine forgery traces through feature reconstruction. The aforementioned fully-supervised temporal forgery localization (FS-TFL) methods have achieved some degree of localization performance. However, they require elaborate frame-level or timestamp annotations for fully-supervised learning, which is usually costly and time-consuming.
@@ 3.4. Deviation Perceiving loss
[…]
@@ 4.1. Experimental Setup
Datasets: We conduct experiments  [^1] The code is available at: https://github.com/wenboxu98/MDP. on two challenging Deepfake temporal partial forgery datasets LAV-DF (Cai et al., 2023) and AV-Deepfake1M (Cai et al., 2024). LAV-DF is a strategic content-driven multimodal forgery dataset, which contains 36,43136,43136 , 431 genuine videos and 99,87399,87399 , 873 forged videos. The duration of forged segments is in the range of $[0-1.6s]$. AV-Deepfake1M is a large-scale multimodal forgery dataset which contains 2,0682,0682 , 068 subjects resulting in 286,721286,721286 , 721 genuine videos and 860,039860,039860 , 039 forged videos. There are four types of samples (real, visual-only forgery, audio-only forgery and audio-visual forgery) in both LAV-DF and AV-Deepfake1M.
@@
Evaluation metrics: The average precision (AP) and average recall (AR) are utilized as the evaluation metrics following (Cai et al., 2023) and (Zhang et al., 2023). For LAV-DF, the IoU thresholds of AP are set as 0.50.50.5, 0.750.750.75 and 0.950.950.95. As AV-Deepfake1M is a more challenging dataset compared to LAV-DF, the IoU thresholds of AP are set as [0.1:0.1:0.7][0.1:0.1:0.7][ 0.1 : 0.1 : 0.7 ]. AR is calculated using 202020, 101010, 555, and 222 proposals with IoU thresholds [0.5:0.05:0.95][0.5:0.05:0.95][ 0.5 : 0.05 : 0.95 ], respectively.
Implementation details: The MDP is trained by Adam optimizer with a learning rate of $1e-5$, a batch size of 32. The hyperparameter $\phi$ is set as 0.50.50.5.
@@ 4.2. Performance Comparisons
LAV-DF Dataset: As shown in Table 1, the results show that MDP, which is a weakly-supervised temporal forgery localization approach, achieves relatively good performance on both AP and AR. Compared to the weakly-supervised temporal action localization approaches, the MDP is significantly improved in both AP and AR. It could be found that the AR@222 to AR@202020 remain consistent, which indicates the proposed framework could predict the forged segments with less number of candidate proposals. As for the AP evaluation metric, MDP achieves significantly superior performance compared to the comparison weakly-supervised approaches on both AP@0.50.50.5 and AP@0.750.750.75, which indicates that MDP exhibits higher accuracy in predicting forged segments. It should be noted that despite MDP showing satisfactory performance on most evaluation metrics, it underperforms on the AP@0.950.950.95, implying that it is inadequate in localizing the precise timestamps of the forged segments.
@@
Obviously, compared to the weakly-supervised approaches, the fully-supervised approaches achieve superior performance on both AP and AR. Such results are also reasonable, as fully-supervised approaches are more adept at learning the relationship between the Deepfake video features and corresponding timestamps of forged segments with provided frame-level annotations. Nevertheless, it could be observed that the MDP also achieves relatively good performance on AP@0.50.50.5 and AP@0.750.750.75 with a relatively small gap with fully-supervised approaches. The experimental results indicate that, despite utilizing only video-level annotations for weakly-supervised learning, the MDP could still effectively identify temporal forgery traces present within the multimodal features, enabling relatively precise localization of timestamps for forged segments. This is achieved by analyzing the temporal deviations in the multimodal features, demonstrating the MDP’s performance in addressing temporal forgery localization.
[…]
@@
The experimental results are shown in Table 2. It could be observed that both fully-supervised approaches and weakly-supervised approach MDP have decreased in AR evaluation metric compared to the experimental results of LAV-DF. As for AP, the proposed MDP underperforms on AP evaluation metrics with large IoU thresholds, while it achieves localization results close to the fully-supervised UMMAFormer on AP@0.10.10.1 and AP@0.20.20.2. The experimental results show that MDP could mine temporal forgery traces and localize the timestamps of forged segments even in challenging AV-Deepfake1M dataset by relying merely on the video-level annotations. We also conducted experiments with WS-TAL approaches CoLA and FuSTAL on the AV-Deepfake1M dataset. The corresponding experimental results were not displayed in Table  2 since the obtained localization results are not effective. The comparison results on this dataset also illustrate that the MDP could exploit the subtle forgery traces within the temporal features, thereby addressing the task of timestamps localization of the forged segments relying on the video-level annotations.
@@
4.3. Ablation Study
[…]
4.4. Visualization Analysis
From the visualization results, it could be observed that MDP could localize the timestamps of all the forged segments more precisely compared to the CoLA and FuSTAL on LAV-DF. Moreover, the localization effectiveness of MDP is comparable to that observed in fully-supervised approaches within the presented sample. In particular, the AV-Deepfake1M sample presents a notable challenge, as the duration of all three forged segments is below 0.30.30.3s, while the overall duration of the Deepfake video is 23.3623.3623.36s. The localization of a small proportion of temporal forgeries is a significant challenge for WS-TFL. Despite this, the MDP effectively utilizes weakly-supervised learning based solely on video-level annotations. It accurately predicts the timestamps of the forged segments, achieving comparable localization results to that obtained through fully-supervised approaches training at the frame-level annotations. This represents that the proposed MDP effectively leverages the temporal forgery traces following the interaction of multimodal features, thereby enabling the precise identification of both the start and end timestamps of the forged segments.